OpenStudy (anonymous):

Find a linear function f(x)=mx+b so that f(f(x))=4x+6. Assume the slop is negative.

OpenStudy (anonymous):

\[ f(f(x))= mf(x) + b= m( mx +b ) +b= m^2 x + m b + b= 4 x + 6 \] Can you finish it?

OpenStudy (anonymous):

\[ m^2 =4\\ m= -2 \text { slope negative} m b + b =6\\ -2 b + b =6\\ -b =6\\ b=-6 \]

OpenStudy (anonymous):

yea i got that far but couldnt see past that please help!

OpenStudy (anonymous):

oooooooooooooohhhhhhhh! Thank you!!!!!

OpenStudy (anonymous):

\[ m^2 =4\\ m= -2 \text { slope negative}\\ m b + b =6\\ -2 b + b =6\\ -b =6\\ b=-6 \]

OpenStudy (anonymous):

yw